OpenStudy (anonymous):

solve the given equation. log27 n= 8/3? show work

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

\[\log_{27}n=\frac{ 8 }{ 3 } \]

OpenStudy (gorv):

\[3*\log_{27}n=8 \]

OpenStudy (gorv):

\[\log_{27}n^{3}=8 \]

OpenStudy (gorv):

\[n^{3}=27^{8}\]

OpenStudy (gorv):

\[n=27^{8*1/3=27^{\frac{ 8 }{ 3 }}=3^{3}}\]

OpenStudy (gorv):



OpenStudy (anonymous):

so the answer is 8/3?? @gorv

OpenStudy (gorv):

3^8=(3*3*3*3*3*3*3*3)

OpenStudy (anonymous):

6,561? @gorv